# Tallahassee, FL 32306-4350, USA

### v1

@@ 13 TeV 36/140 fb-1 CMS:2018lgn /ATLAS:2024ojr
The top-quark interactions have been characterised with excellent precision by the Tevatron and the LHC. Our top-quark data set, summarised in Table 4, contains the most relevant and updated top-quark measurements. This data set consists of differential measurements of top-quark pair production, including also the $t\bar{t}$ charge asymmetry at LHC and the forward-backward asymmetry at Tevatron, differential and inclusive measurements of top-quark production in association with EW boson, [^12] The associated production with Higgs bosons is listed in Section 3.4 and, therefore, not listed here. EW single top-quark production measurements, helicity measurements of top-quark decays, and four heavy quark production measurements including four top-quark production and two top-quark production in association with two bottom quarks. [^13] A subset of these observables have been previously implemented in HEPfit to study specifically the top-quark sector Durieux:2019rbz ; Miralles:2021dyw ; Cornet-Gomez:2025jot .
@@
As in the case of Higgs-boson observables, top-quark observables are also parametrised at linear order in the SMEFT coefficients using MG5_aMC@NLO with the same UFO file, PDF, and scale choices as described in Section 3.4. In this case, however, we have included differential distributions for the $pp\to t\bar{t}$ process with invariant masses reaching up to a few TeV. As shown in ref. Czakon:2016dgf , the most suitable dynamical scale for such differential observables is $H_{T}/4=\frac{1}{4}\left(\sqrt{m_{t}^{2}+p_{T,t}^{2}}+\sqrt{m_{t}^{2}+p_{T,\bar{t}}^{2}}\right),$ which is below one quarter of the invariant mass of the $t\bar{t}$ system. Hence, even for the highest invariant mass bins, the optimal scale choice remains at a few hundred GeV, comparable to the fixed scale $\mu_{W}$ used for the other observables. This suggests that the deviation introduced by using $\mu_{W}$ instead of the optimal dynamic scale should lie within the theoretical uncertainty of our predictions, thereby justifying the choice of keeping $\mu_{W}$ also for this sector.
@@ 3.6 Flavour observables
In the computation of all flavour observables, the SM contribution is evaluated using all available higher-order corrections to achieve the best possible accuracy. The SMEFT Lagrangian is matched to the LEFT at the scale $\mu_{W}$ using the results of ref. Jenkins:2017jig . The LEFT coefficients are then evolved to the relevant energy scale using the LEFT renormalization group equations Jenkins:2017dyc .
@@
$F_{B_{s}}$ (GeV) $0.2301\pm 0.0012$ FLAG:2021npn
$F_{B_{s}}/F_{B_{d}}$ $1.208\pm 0.005$ FLAG:2021npn
$B_{B_{s}}(4.2\,\mathrm{GeV})$ $0.888\pm 0.040$ FLAG:2021npn
$B_{B_{s}}/B_{B_{d}}$ $1.015\pm 0.021$ FLAG:2021npn
$B_{B_{s},4}(4.2\,\mathrm{GeV})$ $0.98\pm 0.08$ FLAG:2021npn
$B_{B_{s},5}(4.2\,\mathrm{GeV})$ $1.66\pm 0.13$ FLAG:2021npn
$B_{B_{d},4}(4.2\,\mathrm{GeV})$ $0.99\pm 0.08$ FLAG:2021npn
$B_{B_{d},5}(4.2\,\mathrm{GeV})$ $1.58\pm 0.18$ FLAG:2021npn
$B_{K}(2\,\mathrm{GeV})$ $0.552\pm 0.012$ FLAG:2021npn
$B_{K,4}(2\,\mathrm{GeV})$ $0.904\pm 0.053$ FLAG:2021npn
$B_{K,5}(2\,\mathrm{GeV})$ $0.618\pm 0.114$ FLAG:2021npn
$\phi_{\varepsilon_{K}}(^{\circ})$ $43.51\pm 0.05$ ParticleDataGroup:2024cfk
$\overline{\kappa}_{\varepsilon_{K}}$ $0.97\pm 0.02$ Ciuchini:2021zgf
$(\Delta M_{K})^{\rm SM}$ (ns-1) $8.8\pm 3.6$ Wang:2022lfq
$B_{D,1}(3\,\mathrm{GeV})$ $0.765\pm 0.025$ FLAG:2021npn
$B_{D,4}(3\,\mathrm{GeV})$ $0.98\pm 0.06$ FLAG:2021npn
$B_{D,5}(3\,\mathrm{GeV})$ $1.05\pm 0.09$ FLAG:2021npn
$F_{D}$ (GeV) $0.2120\pm 0.0007$ FLAG:2021npn
$(\Delta M_{D})^{\rm SM}$ (ps-1) $0.005\pm 0.005$ Betti:2024ldy
$f_{K}$ (GeV) $0.15611\pm 0.00021$ DiCarlo:2019thl
$f_{K}/f_{\pi}$ $1.1966\pm 0.0018$ DiCarlo:2019thl
$\delta R_{\pi}^{\rm phys}$ $0.0153\pm 0.0019$ DiCarlo:2019thl
$\delta P_{c,u}$ $0.04\pm 0.02$ Isidori:2005xm
@@
$\Delta m_{B_{s}}$ $(\mathrm{ps}^{-1})$ $17.765\pm 0.006$ HFLAV:2022esi ; ParticleDataGroup:2024cfk
$\phi_{s}$ $-0.049\pm 0.019$ HFLAV:2022esi ; ParticleDataGroup:2024cfk
$A_{sl}^{s}$ $-0.0006\pm 0.00028$ HFLAV:2022esi ; ParticleDataGroup:2024cfk
$\Delta m_{B_{d}}$ $(\mathrm{ps}^{-1})$ $0.5069\pm 0.0019$ HFLAV:2022esi ; ParticleDataGroup:2024cfk
$S_{J/\psi K_{S}}$ $0.692\pm 0.016$ HFLAV:2022esi ; ParticleDataGroup:2024cfk ; Ciuchini:2005mg
$A_{sl}^{d}$ $-0.0021\pm 0.0017$ HFLAV:2022esi ; ParticleDataGroup:2024cfk
$\Delta M_{K}$ $(\mathrm{ns}^{-1})$ $5.293\pm 0.009$ ParticleDataGroup:2024cfk
$\epsilon_{K}$ $(2.228\pm 0.011)\times 10^{-3}$ ParticleDataGroup:2024cfk
$\phi^{M}_{12}(^{\circ})$ $1.9\pm 1.6$ Betti:2024ldy
${\rm BR}(B\to\tau\nu)\times 10^{4}$ $1.09\pm 0.24$ ParticleDataGroup:2024cfk
${\rm BR}(D\to\tau\nu)\times 10^{4}$ $9.9\pm 1.2$ ParticleDataGroup:2024cfk
${\rm BR}(D\to\mu\nu)\times 10^{4}$ $3.981\pm 0.089$ ParticleDataGroup:2024cfk
${\rm BR}(D_{s}\to\tau\nu)\times 10^{3}$ $5.31\pm 0.11$ ParticleDataGroup:2024cfk
${\rm BR}(D_{s}\to\mu\nu)\times 10^{2}$ $5.37\pm 0.10$ ParticleDataGroup:2024cfk
$\Gamma(K\to\mu\nu)/\Gamma(\pi\to\mu\nu)$ $1.3367\pm 0.0029$ ParticleDataGroup:2024cfk
${\rm BR}(\pi\to\mu\nu)\times 10^{5}$ $3.8408\pm 0.0007$ ParticleDataGroup:2024cfk
$d\Gamma(B\to D\ell\nu)/dw$ $[\Delta\Gamma_{i}/\Delta w]_{10\times 10}$ Belle:2015pkj
${\rm BR}(K^{+}\to\pi^{+}\nu\bar{\nu})\times 10^{10}$ $1.175\pm 0.365$ ParticleDataGroup:2024cfk
${\rm BR}(B\to X_{s}\gamma)\times 10^{4}$ $3.49\pm 0.19$ ParticleDataGroup:2024cfk
$\overline{{\rm BR}}(B_{s}\to\mu^{+}\mu^{-})\times 10^{9}$ $3.41\pm 0.29$ ParticleDataGroup:2024cfk
Table 6: Experimental measurements used for all the flavour observables considered in the present analysis, along with the corresponding references. The set includes both $\Delta F=1$ and $\Delta F=2$ observables, which play a crucial role in constraining the parameters of the Unitarity Triangle analysis UTfit:2022hsi , testing the SM consistency, and probing potential contributions from NP Bona:2024bue .
@@ 3.6.1 $\Delta F=2$ observables
The most general $\Delta F=2$ effective Hamiltonian involves $5+3$ operators Gabbiani:1996hi . However, in the context of the SMEFT, only $3+1$ of them are generated, namely, in the notation of ref. Gabbiani:1996hi :
@@
where $q_{i}=u_{i},d_{i}$ are up- and down-type quark fields, $P_{L,R}=(1\mp\gamma_{5})/2$ are the left- and right-handed projectors, and $\alpha,\beta$ are colour indices. Furthermore, within the $U(2)^{5}$ flavour assumption, the operator $\tilde{Q}_{1}^{q_{i}q_{j}}$ is negligible. In terms of the LEFT operators of ref. Jenkins:2017jig , the Wilson coefficients of the above operators are given by
@@
For the relevant matrix elements we use the lattice B-parameters reported in Table 5. For $B_{q}-\bar{B}_{q}$ mixing ($q=d,s$), we use in the fit the mass difference, the time-dependent CP asymmetries, and the semileptonic asymmetries. For $K-\bar{K}$ mixing, we use the $\epsilon_{K}$ CP-violating parameter and the mass difference $\Delta M_{K}$. For the SM contribution to the latter, we use the lattice QCD estimate in ref. Wang:2022lfq . Finally, for the dispersive phase of $D-\bar{D}$ mixing $\phi^{M}_{12}$, we use the results of the global fit of ref. Betti:2024ldy and, lacking an estimate of the SM contribution to the real part of the amplitude, assume that the SM contribution can saturate the experimental value. All the experimental results used in this section are summarised in Table 6, while the expressions for the observables in terms of the operators in eq. (5) can be found for example in refs. UTfit:2006onp ; UTfit:2007eik .
@@ 3.6.2 $\Delta F=1$ observables
After the RGE below the EW scale, the relevant matrix elements are computed using the lattice QCD results for the decay constants, including the QED corrections for light meson decays DiCarlo:2019thl , as reported in Table 5. The experimental results used in this section are all summarised in Table 6.
For the process $B\to D\ell\nu$, the relevant matching conditions remain those corresponding to the operators shown in eq. (3.6.2). The corresponding matrix elements are computed using the Boyd-Grinstein-Lebed parameterization, employing a $z$-expansion truncated at second order in the conformal variable $z$. In this analysis, we incorporate constraints from lattice QCD given in MILC:2015uhg as well as unitarity bounds on the expansion coefficients. The experimental data used also include a $10\times 10$ correlation matrix ($[\Delta\Gamma_{i}/\Delta w]_{10\times 10}$), see reference in Table 6.
We now turn to the second class of $\Delta F=1$ decays, namely the flavour-changing neutral current (FCNC) processes. In our analysis, we include two key decays that are particularly sensitive to potential contributions from BSM physics: the inclusive radiative decay $B\rightarrow X_{s}\gamma$ and the short-distance dominated exclusive decay $B_{s}\to\mu^{+}\mu^{-}$. SM predictions for these processes are computed to next-to-next-to-leading order accuracy as detailed in references Misiak:2020vlo ; Bobeth:2013uxa . The leading effects of NP for the inclusive radiative channel involve the following operators:
@@
Finally, we consider the FCNC process $K^{+}\to\pi^{+}\nu\bar{\nu}$, which is also theoretically clean and highly suppressed in the SM Brod:2010hi . Due to its strong sensitivity to short-distance physics and minimal hadronic uncertainties Brod:2021hsj , it provides a powerful probe of potential new physics effects. Assuming lepton universality and taking into account the flavour symmetries adopted in our study, the only LEFT operator relevant for this channel is:
@@ 4 Results
In this section, we present the results of the fits we performed with different assumptions on the SMEFT flavour symmetries, namely $U(3)^{5}$ and $U(2)^{5}$ in both the UP and DOWN bases defined in Section 2.1. In order to assess the impact of the RGE of the SMEFT Wilson coefficients, we give results for $\Lambda=10$ and $\Lambda=3$ TeV with RGE, and $\Lambda=1$ TeV without RGE (labeled in the plots and tables as “noRGE”). Given the leading logarithmic accuracy of the RGE, to ensure the dominance of log-enhanced terms we require $\ln{(\mu_{1}^{2}/\mu_{2}^{2})}\gtrsim 2$, and consequently we cannot resolve scale ratios smaller than $\sim 3$, which justifies the chosen values of $\Lambda$. We perform both global fits with all SMEFT Wilson coefficients floating and fits varying one Wilson coefficient at a time (called individual fits in the following), while SM input parameters and hadronic matrix elements are always floated. For individual fits, we also perform fits omitting different sets of observables, to assess the impact of each set on the results. Numeric results for all the fits are reported in Appendix B, while here we illustrate them in a series of plots and discuss their main features and physical interpretation.
@@ 4.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
Assuming that NP respects the $U(3)^{5}$ flavour symmetry, the SMEFT coefficients at the scale $\Lambda$ are given by eq. (8). Let us first consider individual fits. In Table LABEL:tab:u3-ind we report results for $\Lambda=10$ TeV and $\Lambda=3$ TeV including RGE, and for $\Lambda=1$ TeV without RGE. For each coefficient we give the 68% HPDI interval and the bound on $\Lambda/\sqrt{|C_{i}|}$ obtained by taking the maximum of the 95% HPDI interval for $|C_{i}|$. The cases in which the posterior for the relevant Wilson coefficient is not centered around zero so that the $68\%$ or $95\%$ HPDI (from which the bound on $\Lambda/\sqrt{|C_{i}|}$ is obtained) touches the edges of the prior’s range (see Section 3.1) are indicated in red. The $68\%$ HPDI intervals for the $C_{i}$ roughly scale as $\Lambda^{2}$ up to RG effects, unless they hit the prior’s edges. Conversely, the bounds on the effective scale $\Lambda/\sqrt{|C_{i}|}$ only change due to RG effects and possibly due to the prior’s range. It happens for several less constrained coefficients that the sensitivity to experimental bounds gets too diluted at higher scales, so that the $95\%$ HPDI interval for $|C_{i}|$ touches the prior’s edges, and consequently a bound on the scale cannot be obtained. RGE-induced effects are particularly relevant for coefficients such as $C_{G}$, which without RGE is mainly constrained by top-quark observables, [^15] The $C_{G}$ coefficient can also be strongly constrained using multijet production Krauss:2016ely ; Hirschi:2018etq not present in our data set. As shown in those studies, the largest sensitivity to this operator comes from ${\cal O}(1/\Lambda^{4})$ effects, while linear effects provide limited information. or $C_{W}$, mainly constrained by EW observables, while including RGE both get stronger constraints from Higgs-boson observables. This is also illustrated in Fig. 1, based on Table LABEL:tab:u3-ind. In the lower panel, we report the half width of the $95\%$ HPDI interval for $|C_{i}|/\Lambda^{2}$ for the three cases 1) $\Lambda=10$ TeV with RGE, 2) $\Lambda=3$ TeV with RGE, and 3) $\Lambda=1$ TeV without RGE, both in the full fit (i.e. including all observables) and excluding the most constraining set of observables (using the color code explained in the figure). To guide the eye, we also draw the $4\pi/\Lambda^{2}$ perturbativity bound as a horizontal dashed line. The cases in which the $95\%$ HPDI interval touches the prior’s edges, indicated in red in Table LABEL:tab:u3-ind, are hatched with red diagonal lines. When the posterior is flat since the processes considered cannot bound the corresponding coefficient, the $95\%$ HPDI interval is hatched with diagonal white lines. Thus, ranges touching the perturbativity line but solid or hatched in red can be refined by improving the precision of the measurements already included in the fit, while ranges hatched with diagonal white lines can only be improved by including new observables. The results of the fits can also be translated into the allowed range at $95\%$ probability for $\Lambda/\sqrt{|C|}$. Since no coefficient is driven away from zero at $95\%$ probability, we can only put a lower bound on the effective NP scale, as reported in the upper panel of Fig. 1, with the same color code. Finally, we summarise the information on the lower bounds on the effective NP scale obtained for $\Lambda=3$ TeV in Fig. 2.

### v2

@@ 13 TeV 36/140 fb-1 Sirunyan and others (2020b)/Aad and others (2024d)
The top-quark interactions have been characterised with excellent precision by the Tevatron and the LHC. Our top-quark data set, summarised in Table 4, contains the most relevant and updated top-quark measurements. This data set consists of differential measurements of top-quark pair production, including also the $t\bar{t}$ charge asymmetry at LHC and the forward-backward asymmetry at Tevatron, differential and inclusive measurements of top-quark production in association with EW boson, [^13] The associated production with Higgs bosons is listed in Section 3.4 and, therefore, not listed here. EW single top-quark production measurements, helicity measurements of top-quark decays, and four heavy quark production measurements including four top-quark production and two top-quark production in association with two bottom quarks. [^14] A subset of these observables have been previously implemented in HEPfit to study specifically the top-quark sector Durieux et al. (2019); Miralles et al. (2022); Cornet-Gomez et al. (2025).
@@
As in the case of Higgs-boson observables, top-quark observables are also parametrised at linear order in the SMEFT coefficients using MG5_aMC@NLO with the same UFO file, PDF, and scale choices as described in Section 3.4. In this case, however, we have included differential distributions for the $pp\to t\bar{t}$ process with invariant masses reaching up to a few TeV. As shown in ref. Czakon et al. (2017), the most suitable dynamical scale for such differential observables is $H_{T}/4=\frac{1}{4}\left(\sqrt{m_{t}^{2}+p_{T,t}^{2}}+\sqrt{m_{t}^{2}+p_{T,\bar{t}}^{2}}\right),$ which is below one quarter of the invariant mass of the $t\bar{t}$ system. Hence, even for the highest invariant mass bins, the optimal scale choice remains at a few hundred GeV, comparable to the fixed scale $\mu_{W}$ used for the other observables. This suggests that the deviation introduced by using $\mu_{W}$ instead of the optimal dynamic scale should lie within the theoretical uncertainty of our predictions, thereby justifying the choice of keeping $\mu_{W}$ also for this sector.
@@ 3.6 Flavour observables
In the computation of all flavour observables, the SM contribution is evaluated using all available higher-order corrections to achieve the best possible accuracy. The SMEFT Lagrangian is matched to the LEFT at the scale $\mu_{W}$ using the results of ref. Jenkins et al. (2018b). The LEFT coefficients are then evolved to the relevant energy scale using the LEFT renormalization group equations Jenkins et al. (2018a).
@@
$F_{B_{s}}$ (GeV) $0.2301\pm 0.0012$ Aoki and others (2022)
$F_{B_{s}}/F_{B_{d}}$ $1.208\pm 0.005$ Aoki and others (2022)
$B_{B_{s}}(4.2\,\mathrm{GeV})$ $0.888\pm 0.040$ Aoki and others (2022)
$B_{B_{s}}/B_{B_{d}}$ $1.015\pm 0.021$ Aoki and others (2022)
$B_{B_{s},4}(4.2\,\mathrm{GeV})$ $0.98\pm 0.08$ Aoki and others (2022)
$B_{B_{s},5}(4.2\,\mathrm{GeV})$ $1.66\pm 0.13$ Aoki and others (2022)
$B_{B_{d},4}(4.2\,\mathrm{GeV})$ $0.99\pm 0.08$ Aoki and others (2022)
$B_{B_{d},5}(4.2\,\mathrm{GeV})$ $1.58\pm 0.18$ Aoki and others (2022)
$B_{K}(2\,\mathrm{GeV})$ $0.552\pm 0.012$ Aoki and others (2022)
$B_{K,4}(2\,\mathrm{GeV})$ $0.904\pm 0.053$ Aoki and others (2022)
$B_{K,5}(2\,\mathrm{GeV})$ $0.618\pm 0.114$ Aoki and others (2022)
$\phi_{\varepsilon_{K}}(^{\circ})$ $43.51\pm 0.05$ Navas and others (2024)
$\overline{\kappa}_{\varepsilon_{K}}$ $0.97\pm 0.02$ Ciuchini et al. (2022)
$(\Delta M_{K})^{\rm SM}$ (ns-1) $8.8\pm 3.6$ Wang (2022)
$B_{D,1}(3\,\mathrm{GeV})$ $0.765\pm 0.025$ Aoki and others (2022)
$B_{D,4}(3\,\mathrm{GeV})$ $0.98\pm 0.06$ Aoki and others (2022)
$B_{D,5}(3\,\mathrm{GeV})$ $1.05\pm 0.09$ Aoki and others (2022)
$F_{D}$ (GeV) $0.2120\pm 0.0007$ Aoki and others (2022)
$(\Delta M_{D})^{\rm SM}$ (ps-1) $0.005\pm 0.005$ Betti and others (2024)
$f_{K}$ (GeV) $0.15611\pm 0.00021$ Di Carlo et al. (2019)
$f_{K}/f_{\pi}$ $1.1966\pm 0.0018$ Di Carlo et al. (2019)
$\delta R_{\pi}^{\rm phys}$ $0.0153\pm 0.0019$ Di Carlo et al. (2019)
$\delta P_{c,u}$ $0.04\pm 0.02$ Isidori et al. (2005)
@@
$\Delta m_{B_{s}}$ $(\mathrm{ps}^{-1})$ $17.765\pm 0.006$ Amhis and others (2023); Navas and others (2024)
$\phi_{s}$ $0.049\pm 0.019$ Amhis and others (2023); Navas and others (2024)
$A_{sl}^{s}$ $0.0006\pm 0.00028$ Amhis and others (2023); Navas and others (2024)
$\Delta m_{B_{d}}$ $(\mathrm{ps}^{-1})$ $0.5069\pm 0.0019$ Amhis and others (2023); Navas and others (2024)
$S_{J/\psi K_{S}}$ $0.692\pm 0.016$ Amhis and others (2023); Navas and others (2024); Ciuchini et al. (2005)
$A_{sl}^{d}$ $0.0021\pm 0.0017$ Amhis and others (2023); Navas and others (2024)
$\Delta M_{K}$ $(\mathrm{ns}^{-1})$ $5.293\pm 0.009$ Navas and others (2024)
$\epsilon_{K}$ $(2.228\pm 0.011)\times 10^{-3}$ Navas and others (2024)
$\phi^{M}_{12}(^{\circ})$ $1.9\pm 1.6$ Betti and others (2024)
${\rm BR}(B\to\tau\nu)\times 10^{4}$ $1.09\pm 0.24$ Navas and others (2024)
${\rm BR}(D\to\tau\nu)\times 10^{4}$ $9.9\pm 1.2$ Navas and others (2024)
${\rm BR}(D\to\mu\nu)\times 10^{4}$ $3.981\pm 0.089$ Navas and others (2024)
${\rm BR}(D_{s}\to\tau\nu)\times 10^{3}$ $5.31\pm 0.11$ Navas and others (2024)
${\rm BR}(D_{s}\to\mu\nu)\times 10^{2}$ $5.37\pm 0.10$ Navas and others (2024)
$\Gamma(K\to\mu\nu)/\Gamma(\pi\to\mu\nu)$ $1.3367\pm 0.0029$ Navas and others (2024)
${\rm BR}(\pi\to\mu\nu)\times 10^{5}$ $3.8408\pm 0.0007$ Navas and others (2024)
$d\Gamma(B\to D\ell\nu)/dw$ $\Delta\Gamma_{i}/\Delta w]_{10\times 10}$ Glattauer and others (2016)
${\rm BR}(K^{+}\to\pi^{+}\nu\bar{\nu})\times 10^{10}$ $1.175\pm 0.365$ Navas and others (2024)
${\rm BR}(B\to X_{s}\gamma)\times 10^{4}$ $3.49\pm 0.19$ Navas and others (2024)
$\overline{{\rm BR}}(B_{s}\to\mu^{+}\mu^{-})\times 10^{9}$ $3.41\pm 0.29$ Navas and others (2024)
Table 6: Experimental measurements used for all the flavour observables considered in the present analysis, along with the corresponding references. The set includes both $\Delta F=1$ and $\Delta F=2$ observables, which play a crucial role in constraining the parameters of the Unitarity Triangle analysis Bona and others (2023), testing the SM consistency, and probing potential contributions from NP Bona and others (2024).
@@ 3.6.1 $\Delta F=2$ observables
The most general $\Delta F=2$ effective Hamiltonian involves $5+3$ operators Gabbiani et al. (1996). However, in the context of the SMEFT, only $3+1$ of them are generated, namely, in the notation of ref. Gabbiani et al. (1996):
@@
where $q_{i}=u_{i},d_{i}$ are up- and down-type quark fields, $P_{L,R}=(1\mp\gamma_{5})/2$ are the left- and right-handed projectors, and $\alpha,\beta$ are colour indices. Furthermore, within the $U(2)^{5}$ flavour assumption, the operator $\tilde{Q}_{1}^{q_{i}q_{j}}$ is negligible. In terms of the LEFT operators of ref. Jenkins et al. (2018b), the Wilson coefficients of the above operators are given by
@@
For the relevant matrix elements we use the lattice B-parameters reported in Table 5. For $B_{q}-\bar{B}_{q}$ mixing ($q=d,s$), we use in the fit the mass difference, the time-dependent CP asymmetries, and the semileptonic asymmetries. For $K-\bar{K}$ mixing, we use the $\epsilon_{K}$ CP-violating parameter and the mass difference $\Delta M_{K}$. For the SM contribution to the latter, we use the lattice QCD estimate in ref. Wang (2022). Finally, for the dispersive phase of $D-\bar{D}$ mixing $\phi^{M}_{12}$, we use the results of the global fit of ref. Betti and others (2024) and, lacking an estimate of the SM contribution to the real part of the amplitude, assume that the SM contribution can saturate the experimental value. All the experimental results used in this section are summarised in Table 6, while the expressions for the observables in terms of the operators in eq. (5) can be found for example in refs. Bona and others (2006, 2008).
@@ 3.6.2 $\Delta F=1$ observables
After the RGE below the EW scale, the relevant matrix elements are computed using the lattice QCD results for the decay constants, including the QED corrections for light meson decays Di Carlo et al. (2019), as reported in Table 5. The experimental results used in this section are all summarised in Table 6.
For the process $B\to D\ell\nu$, the relevant matching conditions remain those corresponding to the operators shown in eq. (3.6.2). The corresponding matrix elements are computed using the Boyd-Grinstein-Lebed parameterization, employing a $z$-expansion truncated at second order in the conformal variable $z$. In this analysis, we incorporate constraints from lattice QCD given in Bailey and others (2015) as well as unitarity bounds on the expansion coefficients. The experimental data used also include a $10\times 10$ correlation matrix ($\Delta\Gamma_{i}/\Delta w]_{10\times 10}$), see reference in Table 6.
We now turn to the second class of $\Delta F=1$ decays, namely the flavour-changing neutral current (FCNC) processes. In our analysis, we include two key decays that are particularly sensitive to potential contributions from BSM physics: the inclusive radiative decay $B\rightarrow X_{s}\gamma$ and the short-distance dominated exclusive decay $B_{s}\to\mu^{+}\mu^{-}$. SM predictions for these processes are computed to next-to-next-to-leading order accuracy as detailed in references Misiak et al. (2020); Bobeth et al. (2014). The leading effects of NP for the inclusive radiative channel involve the following operators:
@@
Finally, we consider the FCNC process $K^{+}\to\pi^{+}\nu\bar{\nu}$, which is also theoretically clean and highly suppressed in the SM Brod et al. (2011). Due to its strong sensitivity to short-distance physics and minimal hadronic uncertainties Brod et al. (2021), it provides a powerful probe of potential new physics effects. Assuming lepton universality and taking into account the flavour symmetries adopted in our study, the only LEFT operator relevant for this channel is:
@@ 4 Results
[…]
@@ 4.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
Assuming that NP respects the $U(3)^{5}$ flavour symmetry, the SMEFT coefficients at the scale $\Lambda$ are given by eq. (8). Let us first consider individual fits. In Table LABEL:tab:u3-ind we report results for $\Lambda=10$ TeV and $\Lambda=3$ TeV including RGE, and for $\Lambda=1$ TeV without RGE. For each coefficient we give the 68% HPDI interval and the bound on $\Lambda/\sqrt{|C_{i}|}$ obtained by taking the maximum of the 95% HPDI interval for $|C_{i}|$. The cases in which the posterior for the relevant Wilson coefficient is not centered around zero so that the $68\%$ or $95\%$ HPDI (from which the bound on $\Lambda/\sqrt{|C_{i}|}$ is obtained) touches the edges of the prior’s range (see Section 3.1) are indicated in red. The $68\%$ HPDI intervals for the $C_{i}$ roughly scale as $\Lambda^{2}$ up to RG effects, unless they hit the prior’s edges. Conversely, the bounds on the effective scale $\Lambda/\sqrt{|C_{i}|}$ only change due to RG effects and possibly due to the prior’s range. It happens for several less constrained coefficients that the sensitivity to experimental bounds gets too diluted at higher scales, so that the $95\%$ HPDI interval for $|C_{i}|$ touches the prior’s edges, and consequently a bound on the scale cannot be obtained. RGE-induced effects are particularly relevant for coefficients such as $C_{G}$, which without RGE is mainly constrained by top-quark observables, [^16] The $C_{G}$ coefficient can also be strongly constrained using multijet production Krauss et al. (2017); Hirschi et al. (2018) not present in our data set. As shown in those studies, the largest sensitivity to this operator comes from ${\cal O}(1/\Lambda^{4})$ effects, while linear effects provide limited information. or $C_{W}$, mainly constrained by EW observables, while including RGE both get stronger constraints from Higgs-boson observables. This is also illustrated in Fig. 1, based on Table LABEL:tab:u3-ind. In the lower panel, we report the half width of the $95\%$ HPDI interval for $|C_{i}|/\Lambda^{2}$ for the three cases 1) $\Lambda=10$ TeV with RGE, 2) $\Lambda=3$ TeV with RGE, and 3) $\Lambda=1$ TeV without RGE, both in the full fit (i.e. including all observables) and excluding the most constraining set of observables (using the color code explained in the figure). To guide the eye, we also draw the $4\pi/\Lambda^{2}$ perturbativity bound as a horizontal dashed line. The cases in which the $95\%$ HPDI interval touches the prior’s edges, indicated in red in Table LABEL:tab:u3-ind, are hatched with red diagonal lines. When the posterior is flat since the processes considered cannot bound the corresponding coefficient, the $95\%$ HPDI interval is hatched with diagonal white lines. Thus, ranges touching the perturbativity line but solid or hatched in red can be refined by improving the precision of the measurements already included in the fit, while ranges hatched with diagonal white lines can only be improved by including new observables. The results of the fits can also be translated into the allowed range at $95\%$ probability for $\Lambda/\sqrt{|C_{i}|}$. Since no coefficient is driven away from zero at $95\%$ probability, we can only put a lower bound on the effective NP scale, as reported in the upper panel of Fig. 1, with the same color code. Finally, we summarise the information on the lower bounds on the effective NP scale obtained for $\Lambda=3$ TeV in Fig. 2.